OpenStudy (anonymous):

a^4−81 factor completely

OpenStudy (anonymous):

use the following formula \[\Large a^2-b^2=(a+b)(a-b)\]

OpenStudy (lgbasallote):

use difference of two squares \[(a^4 - 81) \implies (a^2)^2 - (9)^2 \implies (a^2 + 9)(a^2 - 9)\] notice how a^2 - 9 is another difference of two squares \[(a^2 + 9)(a^2 - 9) \implies (a^2 + 9)(a^2 - (3^2)^2) \implies (a^2 + 9)(a+3)(a-3)\] does that help?

OpenStudy (lgbasallote):

uhh that's supposed to be \((a^2 - 3^2)\)not (3^2)^2

OpenStudy (lgbasallote):

\[(a^2 + 9)(a^2 - 9) \implies (a^2 + 9)(a^2 + 3^2) \implies (a^2 + 9)(a+3)(a-3)\]

OpenStudy (anonymous):

(a^2+9)(a+3)(a−3) is not correct when i type it in @lgbasallote

OpenStudy (lgbasallote):

that's because there's one more step

OpenStudy (lgbasallote):

\[a^2 + 9 \implies (a+3i)(a-3i)\]

OpenStudy (lgbasallote):

so what do you think will be the final answer?

OpenStudy (anonymous):

(a+3)(a−3)?

OpenStudy (lgbasallote):

no

OpenStudy (lgbasallote):

i told you the answer is \[(a^2 + 9)(a+3)(a-3)\] then i told you \[(a^2 + 9) = (a+3i)(a-3i)\] now try to think what you should do to get the final answer

OpenStudy (anonymous):

(a+3i)(a−3i)(a+3)(a−3)? but where do the "i" come into play

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (lgbasallote):

why there's an i?

OpenStudy (lgbasallote):

and yes that answer is right

OpenStudy (anonymous):

no when i type in (a+3i)(a−3i)(a+3)(a−3) thats not correct

OpenStudy (lgbasallote):

how about (a^2 + 9)(a^2 -9)

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

try doing this one i cant do, 27a^3−8b^3

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

if you just want some answers then go here it's much faster than me http://www.wolframalpha.com/input/?i=27a%5E3%E2%88%928b%5E3